# The Art of Storytelling: Multi-Agent Generative AI for Dynamic Multimodal Narratives

### v3

@@ Abstract
This paper introduces the concept of an education tool that utilizes Generative Artificial Intelligence (GenAI) to enhance storytelling for children. The system combines GenAI-driven narrative co-creation, text-to-speech conversion, and text-to-video generation to produce an engaging experience for learners. We describe the co-creation process, the adaptation of narratives into spoken words using text-to-speech models, and the transformation of these narratives into contextually relevant visuals through text-to-video technology. Our evaluation covers the linguistics of the generated stories, the text-to-speech conversion quality, and the accuracy of the generated visuals.
The Art of Storytelling: Multi-Agent Generative AI for Dynamic Multimodal Narratives
Samee Arif1, Taimoor Arif2, Muhammad Saad Haroon1*, Aamina Jamal Khan1*, Agha Ali Raza1, Awais Athar3, {samee.arif, 25100147, 25100162, agha.ali.raza}@lums.edu.pk taimoora@umich.edu, awais@ebi.ac.uk 1Lahore University of Management Sciences, 2University of Michigan, 3EMBL European Bioinformatics Institute
@@ 1 Introduction
As the role of technology in education grows, there is a unique opportunity to enhance traditional storytelling methods through the use of Generative Artificial Intelligence (GenAI). Recent advancements in Large Language Models (LLMs), Text-to-Speech (TTS), Text-to-Video (TTV) and Text-to-Music (TTM) technologies offer new ways to bring stories to life, creating immersive, multimodal narratives to engage young learners.
In this paper, we introduce a multi-agent system designed to co-create multimodal stories for children. The system is structured around two well-established storytelling frameworks: Freytag’s Pyramid from his book The technique of Drama [^1] https://www.gutenberg.org/files/50616/50616-h/50616-h.htm and Propp’s 31 narrative functions from his book Morphology of the Folktale [^2] https://www.jstor.org/stable/10.7560/783911. Freytag’s Pyramid provides a foundational structure, dividing stories into five phases—exposition, rising action, climax, falling action, and resolution—ensuring that the generated narratives follow a coherent and engaging progression. Propp’s functions serve as a guide for generating the essential elements found in folktales, offering a consistent framework for creating dynamic and imaginative stories that resonate with young audiences.
The system leverages the capabilities of Large Language Models (LLMs) to generate narratives based on user prompts, TTS models to convert these stories into spoken word, TTV models to visualize the narratives through animation, and a TTM model to generate background music that complements the visuals and enhances the storytelling experience. This combination of technologies creates an educational tool, supporting both creative expression and cognitive development, enhancing the overall learning experience for children.
This work presents the use of generative AI in educational applications and also establishes a framework for evaluating the quality and effectiveness of AI-generated content, offering new insights into the role of AI in child-centric educational tools. Our contributions can be summarized as follows:
@@
    System Development: We present a novel multi-agent system that integrates various generative AI models, including TTS models to convert generated stories into emotionally expressive narration, TTV models to generate animated visuals that align with the story structure, and a TTM model to create background music that enhances the emotional tone of the narrative. Together, these components form a cohesive, multimodal storytelling platform that offers a dynamic and immersive experience for children.
@@
    Evaluation Framework: We conduct human evaluation of the system on multiple fronts, including LLMs for story generation and filtering out content inappropriate for children, as well as the TTS and TTV models for their respective roles in the storytelling process. This evaluation ensures that the system delivers content that is both engaging and suitable for young audiences.
@@
    Public Datasets: We make the human evaluation data publicly available, providing a benchmark for assessing LLMs-as-evaluators (Arif et al., 2024) in the story generation task and LLMs-as-content-reviewers. Qwen-VL (Bai et al., 2023) can be tested for video evaluation, and Qwen-Audio (Chu et al., 2023) can be employed for audio evaluation, paving the way for future research in multimodal AI content generation for children. Additionally, we create and release a benchmarking dataset for filtering out content inappropriate for children, further supporting research on child-safe AI systems.
The code, model outputs and the evaluation scores are publicly available on GitHub [^3] https://github.com/ulrs0/The-Art-of-Story-Telling.
@@ 2 System Architecture
Figure 1 shows the multi-agent architecture of our application. The storytelling process is divided into five phases, reflecting Freytag’s Pyramid, with Propp’s narrative functions (e.g., Interdiction, Villainy, and Absentation) distributed across these phases. In each phase, the user selects cards that represent the relevant Propp functions and answers specific questions based on the chosen function. The system front-end design is shown in Appendix A. This approach provides an educational benefit, as it engages children in a structured storytelling process rather than merely asking for direct input, fostering learning about narrative structures and improving their creative thinking.
@@ 3 Related Work
Our paper is centered around four key themes: Generation of stories using LLMs, evaluation of the generated stories, generation and assessment of speech from TTS (text-to-speech) systems, and generation and evaluation of videos from TTV (text-to-video) systems. All these domains have a varying degree of research and prior literature available for us to build our work on.
3.1 Generation of stories using LLMs
The survey paper titled "Automatic Story Generation: A Survey of Approaches" by Alhussain and Azmi (2021) provides a systematic review of various methods in automatic story generation, detailing the evolution from rule-based to AI-driven approaches, and discusses the challenges and future directions in enhancing narrative creativity and coherence. Fan et al. (2018) introduced a method that first generates an outline for a story, followed by the story itself. They used a dataset of user-submitted prompts and user-generated stories for this method. Our methods build on this work by replacing the user who generates the stories with an LLM to improve the speed and scale of the data generated. PlotMachines Rashkin et al. (2020) is a similar system that takes the outline of the story as a set of phrases and generates the story text based on those. The work by Xie and Riedl (2024) takes an iterative approach to story planning. They ask the system a bunch of questions that set up the outline/premise of the story and then generate the story once these details are ironed out. SWAG Patel et al. (2024) is another story-generation framework that uses two models, one that generates the story, and the second that guides the story-generation process. Multiple papers (Zhu et al. (2023), Huang et al. (2023), Ma et al. (2023), Jin et al. (2022)) are specifically targeted toward plot generation and development, which is the focal point of delivering a good story, on top of the structural and linguistic requirements. These methods include step-by-step generation of important parts of the story Zhu et al. (2023), generating and utilizing a sequence of events from a fine-tuned model Ma et al. (2023), and specific tasks using pre-trained models Jin et al. (2022).
@@ 3.2 Story evaluation
There is currently no established gold standard for evaluating stories, particularly those generated by large language models. However, different studies have used their own set of metrics to evaluate the quality of a story from a language, structure, and content perspective. There are quite a few overlapping metrics across these different studies as well. Yang and Jin (2024) list a set of criteria in their survey paper that prior studies have used. These criteria are both story specific e.g. style and language, and LLM specific, e.g. naturalness and non-hallucination. This approach not only measures the quality of the generated text as a story but also checks for common errors associated with LLMs in generation tasks. Akoury et al. (2020) further relied on crowd-sourced evaluation and feedback from user interviews to gauge the quality of their stories—something we aim to do in our future research. Other studies (Chhun et al., 2022) go beyond just the linguistic and structural aspects of the stories but also evaluate the content with metrics like narrative flow and emotions. Researchers from different groups have worked on evaluating the content of the stories from a “story-writing” perspective as well. Xie et al. (2023b) use metrics like fluency, coherence, and Interestingness to present a holistic evaluation framework. Another study by the same group Xie et al. (2023a) evaluates the stories by measuring the closeness between the generated stories and equivalent human-written stories
@@ 4.1 Evaluation
For any system that leverages generative AI for data generation, it is important to rely on a set of evaluation metrics to gauge the quality of the data. Our system uses three forms of data - story text generated through a LLM, audio narration of the stories generated through a TTS model, and the visuals for the stories generated using TTV models. Since each of these data is unique, we need a distinct set of evaluation metrics for each data type. The human evaluators for each module are two Undergraduate in Computer Science students with English as the medium of education. We use weighted Cohen’s Kappa (Cohen, 1960) to find the inter-rater agreement for each evaluation. Each criterion for LLM story generation, TTS, and TTV was evaluated by human raters, with scores assigned on a scale of 0, 1, or 2 (breakdown given in Appendix B).
@@ 4.1.1 LLMs for Story Generation
To evaluate the LLMs listed in Table 1 for story generation, we collected a test dataset. We created a Survey form divided into five parts based on Freytag’s Pyramid. The Propp’s 31 narrative functions were distributed across these sections to align with each part of the narrative structure. Undergraduate students, with English as the medium of education, were asked to fill out the form. In total, we collected 50 test prompts for evaluation.
@@
    Grammar: Is the story grammatically correct? Guan et al. (2019)
@@ 4.1.2 LLMs for Content Moderation
The goal of this evaluation is to ensure that children are not exposed to inappropriate content within stories. To assess the effectiveness of the LLMs as content reviewers, we gathered a diverse dataset of 100 stories—50 appropriate and 50 inappropriate—from Project Gutenberg. These stories were labeled by human annotators based on the presence of violent, explicit, or otherwise unsuitable material for children. This labeled dataset was used to evaluate the LLMs in Table 1. The system prompt for this task is given in Appendix C.
@@ 4.1.4 Text-to-Video Models
For the text-to-video module we evaluate CogVideoX-5b (Yang et al., 2024) on three types of animation styles: cartoon, anime and animated. The ’animated’ style gives the model freehand to generate visuals in any animation style. We sample 50 random paragraphs from the Project Gutenberg free stories and pass them though the Director agent to get the the scene guide. The system prompt for Director is given in Appendix C. The scene guide is then used for video generation. For each video, we assess the following metrics:
@@ 5.1 LLMs for Story Generation
We calculate the weighted Cohen’s Kappa, as shown in table 2, to assess the agreement between total scores by the two raters. The results are mostly above 0.90 indicating strong inter-rater agreement.
Models Kappa ($\kappa$)
GPT-4o-mini 0.93
GPT-4o 0.88
Llama-3.1-8b 0.93
Llama-3.1-70b 0.92
Gemma-2-9b 0.91
Gemma-2-27b 0.92
Table 2: $\kappa$ score between the scoring done by two raters.
Round Test Model Versus Rater 1 Rater 2
Win-Rate Tie-Rate Loss-Rate Win-Rate Tie-Rate Loss-Rate
1 GPT-4o-mini†‡ GPT-4o 60 22 18 54 30 16
Llama-3.1-8b†‡ Llama-3.1-70b 32 46 22 46 32 22
Gemma-2-9b†‡ Gemma-2-27b 36 42 22 34 40 26
2 Llama-3.1-8b†‡ Gemma-2-9b 40 30 30 46 20 34
Gemma-2-9b GPT-4o-mini†‡ 38 18 44 36 18 46
3 Llama-3.1-8b GPT-4o-mini†‡ 34 28 38 32 30 38
@@ 1 GPT-4o-mini†‡ GPT-4o 60 22 18 54 30 16
Table 3: Win-rate, tie-rate, and loss-rate for TTS 1 against TTS 2 based on scoring by both raters. †: Winner according to Rater 1. ‡: Winner according to Rater 2.
To evaluate the LLMs for story generation, we perform a three-round pairwise comparison as shown in Table 3. In Round 1, we compare the models within each family—that is, GPT-4o-mini vs. GPT-4o (OpenAI, 2024); Llama-3.1-8b vs. Llama-3.1-70b (Meta, 2024); and Gemma-2-9b vs. Gemma-2-27b (Google, 2024). Results from Round 1 indicate that within each model family, the smaller models generally outperformed their larger counterparts. Specifically, GPT-4o-mini surpassed GPT-4o with win-rates of 60% and 54%, according to Raters 1 and 2, respectively. Similarly, Llama-3.1-8b achieved win-rates of 32% and 46%, and loss-rates of 22% for both raters against Llama-3.1-70b. Gemma-2-9b attained win-rates of 36% and 34%, and loss-rates of 22% and 26%, respectively, against Gemma-2-27b. These findings suggest that smaller models are more effective for story generation within their respective families.
In Round 2, we compare the winning models from Round 1. Both Rater 1 and 2 prefer Llama-3.1-8b over Gemma-2-9b, with win-rates of 40% and 46%, respectively. In the comparison between Gemma-2-9b and GPT-4o-mini, Gemma-2-9b has a loss-rate of 44% and 46%, with a lower win-rate of 38% and 36%. Although Gemma-2-9b did not emerge as the winner in either comparison, it still demonstrated competitive performance, with a significant portion of ties and relatively close win-rates, indicating the model’s strong performance in generating coherent and engaging stories. In Round 3, the results show that the models are closely matched, with Llama-3.1-8b having win-rates of 34% and 32%, and GPT-4o-mini holding slightly higher win-rates of 38% for both raters. The tie-rates are also notable, with 28% and 30% ties reported by Raters 1 and 2, respectively. The relatively balanced win, loss, and tie rates indicate that both models excel in different aspects of story generation, making the choice between them dependent on specific evaluation criteria.
Figure 2: Comparison of Llama-3.1-8b and GPT-4o-mini for each metric for both raters.
According to Figure 2, GPT-4o-mini has a very slight edge over Llama-3.1-8b in Structural Consistency, Instruction Following and Appropriate Language. However Llama-3.1 has a higher score for Creativity. Given that both models perform so closely, we choose to use Llama-3.1-8b moving forward because it is open-source and has lower computational cost, making it more practical and cost-effective for our needs.

### v4

@@ Abstract
This paper introduces the concept of an education tool that utilizes Generative Artificial Intelligence (GenAI) to enhance storytelling for children. The system combines GenAI-driven narrative co-creation, text-to-speech conversion, text-to-music and text-to-video generation to produce an engaging experience for learners. We describe the co-creation process, the adaptation of narratives into spoken words using text-to-speech models, and the transformation of these narratives into contextually relevant visuals through text-to-video technology. Our evaluation covers the linguistics of the generated stories, the text-to-speech conversion quality, and the accuracy of the generated visuals.
The Art of Storytelling: Multi-Agent Generative AI for
Dynamic Multimodal Narratives
Samee Arif1, Taimoor Arif2, *Muhammad Saad Haroon1, *Aamina Jamal Khan1, Agha Ali Raza1, Awais Athar3 1Lahore University of Management Sciences, 2University of Michigan, 3EMBL European Bioinformatics Institute {samee.arif, 25100147, 25100162, agha.ali.raza}@lums.edu.pk taimoora@umich.edu, awais@ebi.ac.uk
@@ 1 Introduction
In this paper, we introduce a multi-agent system designed to co-create multimodal stories for children. The system is structured around two well-established storytelling frameworks: Freytag’s Pyramid from his book The technique of Drama [^1] https://www.gutenberg.org/files/50616/50616-h/50616-h.htm and Propp’s 31 narrative functions from his book Morphology of the Folktale [^2] https://www.jstor.org/stable/10.7560/783911. Freytag’s Pyramid provides a foundational structure, dividing stories into five phases—exposition, rising action, climax, falling action, and resolution—ensuring that the generated narratives follow a coherent and engaging progression. Propp’s functions serve as a guide for generating the essential elements found in folktales, offering a consistent framework for creating dynamic and imaginative stories that resonate with young audiences. These frameworks can also effectively enhance storytelling skills Ciğerci and Yıldırım (2023). From a pedagogical perspective, these frameworks allow children to actively participate in the storytelling process rather than passively consuming content.
The system leverages the capabilities of Large Language Models (LLMs) to generate narratives based on user prompts, Text-to-Speech (TTS) models to convert these stories into spoken word, Text-to-Video (TTV) models to visualize the narratives through animation, and a Text-to-Music (TTM) model to generate background music that complements the visuals and enhances the storytelling experience. This combination of technologies creates an educational tool, supporting both creative expression and cognitive development, enhancing the overall learning experience for children.
The conversion of written text into spoken words, providing auditory stimuli that can improve reading comprehension, especially for students with learning disabilities. Studies have shown that TTS can support struggling readers by offering audio input as digital text is read aloud, aiding in comprehension and retention (Keelor et al., 2020). Visual stimuli can enhance understanding and retention of information. Research indicates that combining text and images in multimedia explanations is more effective than text alone, as it caters to diverse learning preferences and helps maintain attention (Liu, 2024). Background music adds an emotional layer to narratives, facilitating deeper engagement and supporting memory formation. Music has been recognized for its ability to evoke emotions and create a captivating environment, which can enhance memory retention by increasing engagement (Sinoradzki, 2023). By integrating these components, our system addresses the limitations of traditional storytelling methods that often lack multimodal stimuli, leading to reduced attention spans and engagement. This multimodal approach ensures that storytelling is not only entertaining but also educational, catering to various learning styles and promoting cognitive development.
Our contributions can be summarized as follows:
@@
    System Development: We present a novel multi-agent system that integrates various generative AI models, including TTS models to convert generated stories into emotionally expressive narration, TTV models to generate animated visuals that align with the story structure, and a TTM model to create background music that enhances the emotional tone of the narrative.
@@
    Evaluation Framework: We present an evaluation framework and conduct human evaluation of the system components. Our evaluation includes assessing LLMs for story generation and filtering out content inappropriate for children, as well as the TTS and TTV models for their respective roles in the storytelling process. This evaluation ensures that the system delivers content that is both engaging and suitable for young audiences.
@@
    Public Datasets: We publicly release the human evaluation data, that can be used as a benchmark for assessing LLMs as evaluators in story generation and content review tasks. Additionally, we provide two benchmarking datasets for filtering inappropriate content for children, promoting research on child-safe AI systems.
The code, model outputs and the evaluation scores are publicly available on GitHub [^3] https://github.com/The-Art-of-Story-Telling-2025/The-Art-of-Story-Telling-2025-The-Art-of-Story-Telling-2025.
@@ 2 System Architecture
Figure 1 shows the multi-agent architecture of our application. The storytelling process is divided into five phases, reflecting Freytag’s pyramid, with Propp’s narrative functions (e.g., Interdiction, Villainy and Absentation) distributed throughout these phases. In each phase, the user selects cards that represent the relevant Propp functions and answers specific questions based on the chosen function. The system’s front-end design is shown in Appendix B.
@@ 3 Related Work
3.1 LLMs for Story generation
The survey paper by Alhussain and Azmi (2021) provides a systematic review of various methods in automatic story generation, detailing the evolution from rule-based to AI-driven approaches, and discusses the challenges and future directions in enhancing narrative creativity and coherence. Fan et al. (2018) introduced a method that first generates an outline for a story, followed by the story itself. They used a dataset of user-submitted prompts and user-generated stories for this method. Our methods build on this work by replacing the user who generates the stories with an LLM to improve the speed and scale of the data generated. PlotMachines Rashkin et al. (2020) is a similar system that takes the outline of the story as a set of phrases and generates the story text based on those. The work by Xie and Riedl (2024) takes an iterative approach to story planning. They ask the system a bunch of questions that set up the outline/premise of the story and then generate the story once these details are ironed out. SWAG Patel et al. (2024) is another story-generation framework that uses two models, one that generates the story, and the second that guides the story-generation process. Multiple papers (Zhu et al. (2023), Huang et al. (2023), Ma et al. (2023), Jin et al. (2022)) are specifically targeted toward plot generation and development, which is the focal point of delivering a good story, on top of the structural and linguistic requirements. These methods include step-by-step generation of important parts of the story Zhu et al. (2023), generating and utilizing a sequence of events from a fine-tuned model Ma et al. (2023), and specific tasks using pre-trained models Jin et al. (2022).
@@ 3.2 Story Evaluation
There is currently no established gold standard for evaluating stories, particularly those generated by large language models. However, different studies have used their own set of metrics to evaluate the quality of a story from a language, structure, and content perspective. Yang and Jin (2024) list a set of criteria that prior studies have used. These criteria are both story specific e.g. style and language, and LLM specific, e.g. naturalness and non-hallucination. This approach not only measures the quality of the generated text as a story but also checks for common errors associated with LLMs in generation tasks. Akoury et al. (2020) further relied on crowd-sourced evaluation and feedback from user interviews to gauge the quality of their stories. Other studies (Chhun et al., 2022) go beyond just the linguistic and structural aspects of the stories but also evaluate the content with metrics like narrative flow and emotions. Researchers from different groups have worked on evaluating the content of the stories from a “story-writing” perspective as well. Xie et al. (2023b) use metrics like fluency, coherence, and Interestingness to present a holistic evaluation framework. Another study by the same group Xie et al. (2023a) evaluates the stories by measuring the closeness between the generated stories and equivalent human-written stories
@@ 4.1 Evaluation
For any system that leverages generative AI for data generation, it is important to rely on a set of evaluation metrics to gauge the quality of the data. Our system uses three forms of data - story text generated through a LLM, audio narration of the stories generated through a TTS model, and the visuals for the stories generated using TTV models. Since each of these data is unique, we need a distinct set of evaluation metrics for each data type. The human evaluators for each module are six Computer Science students with English as the medium of education. Each criterion for LLM story generation, TTS, and TTV was evaluated by human raters, with scores assigned on a scale of 0, 1, or 2 (breakdown given in Appendix C).
@@ 4.1.1 LLMs for Story Generation
To evaluate the LLMs listed in Table 1 for story generation, we collected a test dataset. We created a Survey form divided into five parts based on Freytag’s Pyramid. Propp’s 31 narrative functions were distributed across these sections to align with each part of the narrative structure. Undergraduate students, with English as the medium of education, were asked to fill out the form. In total, we collected 50 test prompts for evaluation.
@@
    Grammar: Is the story grammatically correct? (Guan et al., 2019)
@@ 4.1.2 LLMs for Content Moderation
The goal of this evaluation is to ensure that children are not exposed to inappropriate content within stories. To assess the effectiveness of the LLMs as content reviewers, We gathered a diverse dataset of 100 stories from Projecgt Gutenberg, 50 appropriate and 50 inappropriate as rated by our evaluators.. Additionally, we generated a dataset of 50 stories using LLMs, which were similarly labeled as appropriate or inappropriate by human annotators. These stories were labeled by human annotators based on the presence of violent, explicit, or otherwise unsuitable material for children. This labeled dataset was used to evaluate the LLMs in Table 1. The system prompt for this task is given in Appendix D.
@@ 4.1.4 Text-to-Video Models
For the TTV module we evaluate CogVideoX-5b (Yang et al., 2024) on three types of animation styles: cartoon, anime and animated. The ‘animated‘ style gives the model freehand to generate visuals in any animation style. We sample 50 random paragraphs from the Project Gutenberg free stories and pass them though the Director agent to get the the scene guide. The system prompt for Director is given in Appendix D. The scene guide is then used for video generation. For each video, we assess the following metrics:
@@ 5 Results and Discussion
Throughout this section, the win rate reflects the proportion of times a model outperformed others, the tie rate indicates instances of comparable performance, and the loss rate captures the frequency of underperformance.
@@ 5.1 LLMs for Story Generation
Table 2 summarizes the pairwise comparisons among six models of varying scales—Gemma-2-9b, Gemma-2-27b (Google, 2024), Llama-3.1-8b, Llama-3.1-70b (Meta, 2024), GPT-4o, and GPT-4o-mini (OpenAI, 2024)—based on human evaluations of their story outputs.
A first notable observation is that simply scaling up parameter counts does not guarantee stronger performance in the downstream task of story generation. Gemma-2-9b outperforms its larger counterpart Gemma-2-27b (39.67% vs. 25.67% in win-rate), underscoring that bigger parameter counts are not always decisive for story-generation quality. Likewise Llama-3.1-70b achieves notable 43.33% wins and only 27.67% losses against GPT-4o. GPT-4o also performs poorly against GPT-4o-mini with only 16.33% win-rate. Across most pairwise comparisons, tie-rates hover near 30%, suggesting that many of these models are closely matched in generating coherent and engaging stories. Based on these head-to-head outcomes, Gemma-2-9b emerges as the most consistently strong performer overall, outperforming all the other models. Interestingly, the bigger model from the same family, Gemma-2-27b is the weakest based on the human evaluation, wining none of the pairwise comparisons.
Rates (%)
Test Model Versus Model Win Tie Loss
Gemma-9b Gemma-27b 39.67 34.67 25.67
Gemma-9b Llama-8b 39.17 34.00 26.83
Gemma-9b Llama-70b 37.44 35.33 27.22
Gemma-9b GPT-4o 40.00 33.42 26.58
Gemma-9b GPT-4o-mini 38.27 33.60 28.13
Gemma-27b Llama-8b 31.33 32.33 36.33
Gemma-27b Llama-70b 28.33 36.00 35.67
Gemma-27b GPT-4o 31.56 33.67 34.78
Gemma-27b GPT-4o-mini 29.92 33.08 37.00
Llama-8b Llama-70b 28.33 41.00 30.67
Llama-8b GPT-4o 34.00 38.17 27.83
Llama-8b GPT-4o-mini 30.33 37.78 31.89
Llama-70b GPT-4o 43.33 29.00 27.67
Llama-70b GPT-4o-mini 34.00 33.67 32.33
GPT-4o GPT-4o-mini 16.33 36.67 47.00
@@
Table 2: Win-rate, tie-rate, and loss-rate for Test Model against Versus Model based on scoring by all raters.
Figure 2: Comparison of the six LLMs for each metric based on average human-evaluation scores.
Figure 2 presents average human-evaluation scores (on a 0–2 scale) across seven criteria—Grammar, Linguistic Consistency, Appropriate Language, Structural Consistency, Creativity, Adherence to Instructions, and Naturalness. A closer look at Grammar scores shows that GPT-4o-mini leads with an average of 1.96, while most other models hover around 1.80–1.90. For Linguistic Consistency, GPT-4o and GPT-4o-mini both top out at 1.90, whereas Llama-3.1-8b, Llama-3.1-70b, and Gemma-2-9b trail slightly in the 1.83–1.88 range, suggesting all models are fairly consistent in using language of similar register or difficulty throughout a story. In Appropriate Language, GPT-4o-mini again scores highest (1.97), closely followed by the Gemma-2 series at 1.94, signaling strong ability to avoid potentially offensive or overly complex wording.
In Structural Consistency, Gemma-2-9b (1.81) stands out as the most coherent, edging out Gemma-2-27b, GPT-4o-mini, and Llama-3.1-70b (all near 1.78) while GPT-4o (1.66) lags notably. In Creativity, Gemma-2-9b achieves the top mean score (1.65), surpassing all other models—especially GPT-4o at 1.37—indicating Gemma-2-9b’s stronger capacity for generating engaging stories. Under Adherence to Instructions, Llama-3.1-8b and GPT-4o-mini both earn the highest score (1.93), followed closely by Gemma-2-9b (1.92) and Llama-3.1-70b (1.91), showcasing that all models tend to follow user prompts precisely. Lastly, Naturalness is also led by GPT-4o-mini at 1.87, with Llama-3.1-8b just behind (1.86), reflecting how “human-like” or fluid their story outputs appear to human readers.
Overall, GPT-4o-mini emerges as most consistent across the board, consistently appearing at or near the top in most categories, especially Grammar, Appropriate Language, and Naturalness. Gemma-2-9b scores highest in Creativity and Structural Consistency, making it appealing for users. In contrast, GPT-4o sits at the lower end in several metrics (particularly Creativity and Structural Consistency), though it remains reasonably competitive in Grammar and Linguistic Consistency.